two years ago kumar a father was 5 times as old as a son.2 years later,his age will be 8 years more than 3 times the age of his son.find the present age of father and son
Answer by richwmiller(17219)

f-2=5*(s-2);
does two years later mean from the 2 years before or two years from now
two years from now
f+2=8+3*(s+2)
f=42 s=10
two years from before ie now
f=8+3*(s)
f=32 s=8